Imagine that you performed this experiment as described in your lab manual. In doing so, your first temperature reading was 32℃ , and your second reading was 54℃. Your first volume was 21 mL . what would be your calculated 2nd volume?
Answer #1

According to Boyle's law:

V1/T1 = V2/T2

Where V1 and V2 are the first and second volumes.

T1 and T2 are the first and second absolute temperatures (in kelvin)

i.e. 21 mL/(32+273) K = V2/(54+273) K

Therefore, the second volume (V2) = 22.5 mL